# Automated Proof Generation for Rust Code via Self-Evolution

### v2

@@ Abstract
Ensuring correctness is crucial for code generation. Formal verification offers a definitive assurance of correctness, but demands substantial human effort in proof construction and hence raises a pressing need for automation. The primary obstacle lies in the severe lack of data—there is much fewer proofs than code snippets for Large Language Models (LLMs) to train upon. In this paper, we introduce SAFE, a framework that overcomes the lack of human-written proofs to enable automated proof generation of Rust code. SAFE establishes a self-evolving cycle where data synthesis and fine-tuning collaborate to enhance the model capability, leveraging the definitive power of a symbolic verifier in telling correct proofs from incorrect ones. SAFE also re-purposes the large number of synthesized incorrect proofs to train the self-debugging capability of the fine-tuned models, empowering them to fix incorrect proofs based on the verifier’s feedback. SAFE demonstrates superior efficiency and precision compared to GPT-4o. Through tens of thousands of synthesized proofs and the self-debugging mechanism, we improve the capability of open-source models, initially unacquainted with formal verification, to automatically write proofs for Rust code. This advancement leads to a significant improvement in performance, achieving a 52.52% accuracy rate in a benchmark crafted by human experts, a significant leap over GPT-4o’s performance of 14.39%.
@@ 1 Introduction
Next, to obtain tens of thousands of verified Rust programs, SAFE similarly uses a self-evolving procedure. Code proofs are very difficult to synthesize: even with careful prompt engineering, GPT-4o managed to synthesize correct proofs for only $<20\%$ of programs in the the preceding dataset with synthesized specifications, after a month of non-stop invocations. Fortunately, this is sufficient to bootstrap SAFE. Through rounds of self-evolving, the quantity and quality of synthesized proofs keep increasing, while fine-tuned open-source models’ capability of proof synthesis keeps getting augmented. These models also generate proofs much faster than GPT-4o. The keys to the success here include (1) the ability of Verus in authoritatively and quickly telling correct proofs from incorrect ones, allowing SAFE to sift through the huge amount of low-quality data in early rounds without any human labeling, and (2) the reasonable quality of SAFE-synthesized specifications that allow SAFE to largely avoid trivial proofs, which have little usage in fine-tuning.
@@ 2.1 Verus verification tool
Verus (Lattuada et al., 2023) statically analyzes every proof target (i.e., a Rust function and its specification) and any given proof annotations, and forms queries for the underlying SMT solver (e.g., Z3 (De Moura & Bjørner, 2008)) to solve. By leveraging the type system in Rust and allowing both specification and proof annotations to be written in Rust syntax, Verus has become the state-of-the-art verification tool for Rust, one of the most popular programming languages (Perkel, 2020; Fulton et al., 2021), and has been used to successfully verify large systems (Sun et al., 2024b; Zhou et al., 2024). Unfortunately, since Verus has been developed for only about three years, there are fewer than 500 Verus verified Rust files on GitHub.
@@ 3 Approach
As illustrated in Figure 1, SAFE involves two self-evolving (Tao et al., 2024; Gulcehre et al., 2023) procedures. The first procedure synthesizes specifications that are used as inputs to the second procedure, while the second procedure produces the end-result of SAFE—a fine-tuned LLM that can automatically synthesize proofs for Rust code. As listed in Algorithm 1, both procedures use GPT-4o to generate the round-00 data $spec/proof\textit{-}data_{0}$. In each round $r$, we use data collected and filtered from earlier rounds, $10...r-$ 1, to fine-tune $\text{model}_{r}$ based on its preceding $\text{model}_{r-1}$. At the end of each round, $\text{model}_{r}$ is used to generate the round-$r$ data. High-quality specifications are preserved for the next round. Low-quality specifications are discarded; incorrect proofs are re-purposed to become self-debugging training data for the next round.
@@
1
Input : $model_{0}$, an open-source LLM; $rust\_programs$, a set of Rust programs
@@
2 $programs\leftarrow\mbox{GPT-4o}.translate(rust\_programs).compilable(\mbox{Verus Compiler})$;
3 Specification Generation :
4       $spec\textit{-}data_{0}\leftarrow\mbox{GPT-4o}.generate(programs).filter(Tests)$;
5       for $r\mbox{ in }[1,\dots R_{1}]$ do
6             $model_{r}\leftarrow model_{r-1}.fine\textit{-}tune(\bigcup\limits_{i=1}^{r-1}spec\textit{-}data_{i})$;
7             $spec\textit{-}data_{r}\leftarrow model_{r}.generate(programs).filter(Tests)$;
8
9      return $model_{r}$;
10
11
12Auto-Proving :
13       $proof\textit{-}data_{0},debug\textit{-}data_{0}\leftarrow\mbox{GPT-4o}.generate(programs).filter(\mbox{Verus Verifier})$;
14       for $r\mbox{ in }[1,\dots R_{2}]$ do
15             $model_{r}\leftarrow model_{r-1}.fine\textit{-}tune(\bigcup\limits_{i=1}^{r-1}(proof\textit{-}data_{i}+debug\textit{-}data_{i}))$;
16             $proof\textit{-}data_{r}\leftarrow model_{r}.generate(programs).filter(\mbox{Verus Verifier})$;
17             $debug\textit{-}data_{r}\leftarrow model_{r}.debug(programs).filter(\mbox{VerusVerifier})$;
18
19      return $model_{r}$;
20
@@ 3.2 Step 2: Self-evolving Specification Synthesis
In this step, LLMs are tasked with generating preconditions and postconditions for a Rust function based on the function’s implementation and docstring. Different from prior work in specification generation (Flanagan & Leino, 2001; Ma et al., 2024), SAFE needs a synthesized specification as input for its self-evolving proof-synthesis framework, which raises unique requirements on its specification evaluation criteria (i.e., which specification to keep or discard) and the mechanism to evaluate the criteria.
@@
We can now set a threshold that fits the need of SAFE: we keep all the specifications with $\geq$80Correctness score and $\geq$60of Completeness score. This way, SAFE can obtain many imperfect but useful specifications for its proof synthesis. We do not set the Correctness threshold to be 100%, because we use GPT to generate test cases when few test cases are provided by the dataset, and we cannot guarantee perfect correctness of these test cases. Since Verus can symbolically evaluate $S(i,o)$ in the preceding formula without checking or running the Rust function, the scoring is fast.
@@ 3.3 Step 3: Self-evolving Proof Synthesis
$\vspace{-0.2cm}\mathcal{L}_{Gen}(\theta)=-\sum_{i=0}^{N}\log P_{\theta}(y_{i}\mid S,y_{t<i})\vspace{-0.2cm}$ (2)
@@ 4.3 Evaluation Results
[…]
Table 1 also shows the efficacy of self-debugging (i.e., SAFE+). For example, after one initial proof is sampled, applying self-debugging on this proof to produce the second sample (Accuracy@2 for SAFE+) consistently outperforms simply producing a second sample of proof without debugging (Accuracy@2 for SAFE) for both LLaMa and DeepSeekCoder on both VerusBench and CodeNet-Test, as shown in Table 1. SAFE+ with DeepSeekCoder achieves the best performance of 70.50% in Table 1 when 10$\times$10 debugged proofs are produced for 10 initial proofs for each proof task.
@@
Self- Evolving Metric Proof Generation Self-Debugging ($K+K*K$)
@@ 4.3.3 Impact of Specification Quality
As mentioned in Section 3.2, we keep only specifications with high Correctness scores and reasonably high Completeness scores. To investigate the impact of this design decision, we zoom into the last round of proof-synthesis model fine-tuning to see how the quality of a proof-synthesis model can be affected by the quality of the used specification dataset. Specifically, the current Round-3 SAFE model is obtained by fine-tuning the Round-2 SAFE model using correct proofs synthesized by the bootstrapping model, the Round-1 model, and the Round-2 model based on the specifications $S$ selected by SAFE, denoted as $\mathbf{P_{0}[\mathbf{S}]}\cup\mathbf{P_{1}[\mathbf{S}]}\cup\mathbf{P_{2}[\mathbf{S}]}$ ($\mathbf{P}_{i}[\mathbf{S}]$ denotes all the correct proofs synthesized in round $i$ based on specification set $S$). In our two alternative settings, we go back to the set of all specifications synthesized during the specification synthesis self-evolving procedure and randomly sample a set of specifications $S*$ that have the same size as $S$; naturally $S*$ contains many low-quality specifications that do not pass our Correctness or Completeness threshold. In our alternative setting of ‘Mix-Quality’, we obtain a model by fine-tuning the Round-2 SAFE model using $\mathbf{P_{0}[\mathbf{S}]}\cup\mathbf{P_{1}[\mathbf{S}]}\cup\mathbf{P_{2}[\mathbf{S*}]}$; in our alternative setting of ‘Low-Quality’, we obtain a model by fine-tuning the original DeepSeekCoder using $\mathbf{P_{2}[\mathbf{S*}]}$ alone.
@@
Model Metric Proof Generation Self-Debugging ($K+K*K$)
@@ C.1 Specification Filtering
As mentioned earlier in the paper, to evaluate the completeness of a specification, we mutate existing test cases to see whether incorrect test cases can be rejected by a specification. In this example, we mutate the output in the ground truth by adding a new value $15$ into $result$ vector. In this case, Verus will report that the postcondition on Line 7 fails because $15$ does not exist in vector $a$. If a synthesized specification does not contain something like Line 7, this mutated test case will likely point out the incompleteness of that specification.
@@ E.2 Distribution of Specification Metrics
Figure 4 shows the distribution of the correctness-score and the completeness-score of all the specifications synthesized during the self-evolving process of SAFE. As we can see, during the self-evolving process, many trivially correct specifications are generated (i.e., $1.0$ correctness score but $<0.6$ completeness score)—the clusters at the top of the figure and not overlapping with the gray rectangle. These specifications tend to lead to trivial proofs and hence are filtered out. There is also a big cluster of incorrect specifications that have incorrectly rejected all test cases at the bottom-right corner of the figure (i.e., 00 correctness score but $1.0$ completeness score). There is no way Verus can prove that the Rust function satisfies these specifications, no matter what proof annotations are used, so it is good that SAFE filters all of them out.
@@ E.4 Self-Evolution with Small Backbone Models
Proof Generation Self Debugging ($k+k*k$)
@@ F.2 The Effectiveness of Our Self-Debugging Mechanism
Listing 10: An incorrect proof of $is\_sorted$ generated by SAFE before self-debugging.
@@
Listing 11: The correct proof of $is\_sorted$ fixed by SAFE after self-debugging.
@@ F.3 An Ineffective Example
Listing 12: An incorrect proof of $remove$_$odds$ before and after self-debugging.
@@
Listing 13: The ground truth proof of $remove\_odds$.
@@ F.4 An Example Proof with a Low-Quality Specification
Listing 14: An example proof of $gcd$ with a low quality specification.

### v3

@@ Abstract
Ensuring correctness is crucial for code generation. Formal verification offers a definitive assurance of correctness, but demands substantial human effort in proof construction and hence raises a pressing need for automation. The primary obstacle lies in the severe lack of data—there is much fewer proofs than code snippets for Large Language Models (LLMs) to train upon. In this paper, we introduce SAFE, a framework that overcomes the lack of human-written proofs to enable automated proof generation of Rust code. SAFE establishes a self-evolving cycle where data synthesis and fine-tuning collaborate to enhance the model capability, leveraging the definitive power of a symbolic verifier in telling correct proofs from incorrect ones. SAFE also re-purposes the large number of synthesized incorrect proofs to train the self-debugging capability of the fine-tuned models, empowering them to fix incorrect proofs based on the verifier’s feedback. SAFE demonstrates superior efficiency and precision compared to GPT-4o. Through tens of thousands of synthesized proofs and the self-debugging mechanism, we improve the capability of open-source models, initially unacquainted with formal verification, to automatically write proofs for Rust code. This advancement leads to a significant improvement in performance, achieving a 52.52% accuracy rate in a benchmark crafted by human experts, a significant leap over GPT-4o’s performance of 14.39%. The training dataset produced by this work is available on HuggingFace [^1] https://huggingface.co/datasets/microsoft/Verus_Training_Data.
@@ 1 Introduction
Next, to obtain tens of thousands of verified Rust programs, SAFE similarly uses a self-evolving procedure. Code proofs are very difficult to synthesize: even with careful prompt engineering, GPT-4o managed to synthesize correct proofs for only $20\%$ of programs in the the preceding dataset with synthesized specifications, after a month of non-stop invocations. Fortunately, this is sufficient to bootstrap SAFE. Through rounds of self-evolving, the quantity and quality of synthesized proofs keep increasing, while fine-tuned open-source models’ capability of proof synthesis keeps getting augmented. These models also generate proofs much faster than GPT-4o. The keys to the success here include (1) the ability of Verus in authoritatively and quickly telling correct proofs from incorrect ones, allowing SAFE to sift through the huge amount of low-quality data in early rounds without any human labeling, and (2) the reasonable quality of SAFE-synthesized specifications that allow SAFE to largely avoid trivial proofs, which have little usage in fine-tuning.
@@ 2.1 Verus verification tool
Verus (Lattuada et al., 2023) statically analyzes every proof target (i.e., a Rust function and its specification) and any given proof annotations, and forms queries for the underlying SMT solver (e.g., Z3 (De Moura and Bjørner, 2008)) to solve. By leveraging the type system in Rust and allowing both specification and proof annotations to be written in Rust syntax, Verus has become the state-of-the-art verification tool for Rust, one of the most popular programming languages (Perkel, 2020; Fulton et al., 2021), and has been used to successfully verify large systems (Sun et al., 2024b; Zhou et al., 2024). Unfortunately, since Verus has been developed for only about three years, there are fewer than 500 Verus verified Rust files on GitHub.
@@ 3 Approach
As illustrated in Figure 1, SAFE involves two self-evolving (Tao et al., 2024; Gulcehre et al., 2023) procedures. The first procedure synthesizes specifications that are used as inputs to the second procedure, while the second procedure produces the end-result of SAFE—a fine-tuned LLM that can automatically synthesize proofs for Rust code. As listed in Algorithm 1, both procedures use GPT-4o to generate the round-0 data $spec/proof\textit{-}data_{0}$. In each round $r$, we use data collected and filtered from earlier rounds, 0r10...r-1, to fine-tune $\text{model}_{r}$ based on its preceding $\text{model}_{r-1}$. At the end of each round, $\text{model}_{r}$ is used to generate the round-$r$ data. High-quality specifications are preserved for the next round. Low-quality specifications are discarded; incorrect proofs are re-purposed to become self-debugging training data for the next round.
@@
Input : $model_{0}$, an open-source LLM; rust_programsrust\_programs, a set of Rust programs
@@
$programs\leftarrow\mbox{GPT-4o}.translate(rust\_programs).compilable(\mbox{Verus Compiler})$;
Specification Generation :
 $spec\textit{-}data_{0}\leftarrow\mbox{GPT-4o}.generate(programs).filter(Tests)$;
 for $r\mbox{ in }[1,\dots R_{1}]$ do
    $model_{r}\leftarrow model_{r-1}.fine\textit{-}tune(\bigcup\limits_{i=1}^{r-1}spec\textit{-}data_{i})$;
    $spec\textit{-}data_{r}\leftarrow model_{r}.generate(programs).filter(Tests)$;
   end for
 return $model_{r}$;
Auto-Proving :
 $proof\textit{-}data_{0},debug\textit{-}data_{0}\leftarrow\mbox{GPT-4o}.generate(programs).filter(\mbox{Verus Verifier})$;
 for $r\mbox{ in }[1,\dots R_{2}]$ do
    $model_{r}\leftarrow model_{r-1}.fine\textit{-}tune(\bigcup\limits_{i=1}^{r-1}(proof\textit{-}data_{i}+debug\textit{-}data_{i}))$;
    $proof\textit{-}data_{r}\leftarrow model_{r}.generate(programs).filter(\mbox{Verus Verifier})$;
    $debug\textit{-}data_{r}\leftarrow model_{r}.debug(programs).filter(\mbox{Verus Verifier})$;
   end for
 return $model_{r}$;
@@ 3.2 Step 2: Self-evolving Specification Synthesis
In this step, LLMs are tasked with generating preconditions and postconditions for a Rust function based on the function’s implementation and docstring. Different from prior work in specification generation (Flanagan and Leino, 2001; Ma et al., 2024), SAFE needs a synthesized specification as input for its self-evolving proof-synthesis framework, which raises unique requirements on its specification evaluation criteria (i.e., which specification to keep or discard) and the mechanism to evaluate the criteria.
@@
We can now set a threshold that fits the need of SAFE: we keep all the specifications with $\geq$80% Correctness score and $\geq$60% of Completeness score. This way, SAFE can obtain many imperfect but useful specifications for its proof synthesis. We do not set the Correctness threshold to be 100%, because we use GPT to generate test cases when few test cases are provided by the dataset, and we cannot guarantee perfect correctness of these test cases. Since Verus can symbolically evaluate $S(i,o)$ in the preceding formula without checking or running the Rust function, the scoring is fast.
@@ 3.3 Step 3: Self-evolving Proof Synthesis
$\vskip-5.69046pt\mathcal{L}_{Gen}(\theta)=-\sum_{i=0}^{N}\log P_{\theta}(y_{i}\mid S,y_{t<i})\vskip-5.69046pt$ (2)
@@ 4.3 Evaluation Results
We can see that SAFE achieves substantially higher accuracy compared to baseline approaches in both datasets and metrics. In the VerusBench benchmark, even without self-debugging, SAFE achieves an Accuracy@1 of 46.04% and an Accuracy@10 of 53.96% while the best of ‘Raw’ results come from GPT-4o with 11.51% in Accuracy@1 and 24.46% in Accuracy@10. With the long and carefully designed prompt, which was used to bootstrap SAFE proof synthesis, GPT-4o is reasonably effective for VerusBench. For the CodeNet-Test benchmark, we find the performance of baselines and prompt-based GPT-4o, denoted as Prompt, drops significantly—by more than 10XX compared with VerusBench, while SAFE does not suffer as much. It is important to note that CodeNet-Test is significantly larger and exhibits a distinct data distribution compared to VerusBench, and the former is derived from competitive programming, whereas the latter encompasses common algorithms. The results indicate that the prompt-based approach has limited generalizability in automated proof generation, while SAFE no longer needs complicated prompts during inference. We further report paired t-test (Hsu and Lachenbruch, 2014) results for Table 1 in Table 5 in Appendix.
Table 1 also shows the efficacy of self-debugging (i.e., SAFE+). For example, after one initial proof is sampled, applying self-debugging on this proof to produce the second sample (Accuracy@2 for SAFE+) consistently outperforms simply producing a second sample of proof without debugging (Accuracy@2 for SAFE) for both LLaMa and DeepSeekCoder on both VerusBench and CodeNet-Test, as shown in Table 1. SAFE+ with DeepSeekCoder achieves the best performance of 70.50% in Table 1 when $\times10$ debugged proofs are produced for 10 initial proofs for each proof task.
@@
Self- Evolving Metric Proof Generation Self-Debugging (K+KKK+K*K)
@@ 4.3.3 Impact of Specification Quality
As mentioned in Section 3.2, we keep only specifications with high Correctness scores and reasonably high Completeness scores. To investigate the impact of this design decision, we zoom into the last round of proof-synthesis model fine-tuning to see how the quality of a proof-synthesis model can be affected by the quality of the used specification dataset. Specifically, the current Round-3 SAFE model is obtained by fine-tuning the Round-2 SAFE model using correct proofs synthesized by the bootstrapping model, the Round-1 model, and the Round-2 model based on the specifications $S$ selected by SAFE, denoted as $\mathbf{P_{0}[\mathbf{S}]}\cup\mathbf{P_{1}[\mathbf{S}]}\cup\mathbf{P_{2}[\mathbf{S}]}$ ($\mathbf{P}_{i}[\mathbf{S}]$ denotes all the correct proofs synthesized in round $i$ based on specification set $S$). In our two alternative settings, we go back to the set of all specifications synthesized during the specification synthesis self-evolving procedure and randomly sample a set of specifications SS* that have the same size as $S$; naturally SS* contains many low-quality specifications that do not pass our Correctness or Completeness threshold. In our alternative setting of ‘Mix-Quality’, we obtain a model by fine-tuning the Round-2 SAFE model using $\mathbf{P_{0}[\mathbf{S}]}\cup\mathbf{P_{1}[\mathbf{S}]}\cup\mathbf{P_{2}[\mathbf{S*}]}$; in our alternative setting of ‘Low-Quality’, we obtain a model by fine-tuning the original DeepSeekCoder using $\mathbf{P_{2}[\mathbf{S*}]}$ alone.
@@
Model Metric Proof Generation Self-Debugging (K+KKK+K*K)
@@ C.1 Specification Filtering
As mentioned earlier in the paper, to evaluate the completeness of a specification, we mutate existing test cases to see whether incorrect test cases can be rejected by a specification. In this example, we mutate the output in the ground truth by adding a new value $15$ into resultresult vector. In this case, Verus will report that the postcondition on Line 7 fails because $15$ does not exist in vector $a$. If a synthesized specification does not contain something like Line 7, this mutated test case will likely point out the incompleteness of that specification.
@@ E.2 Distribution of Specification Metrics
Figure 4 shows the distribution of the correctness-score and the completeness-score of all the specifications synthesized during the self-evolving process of SAFE. As we can see, during the self-evolving process, many trivially correct specifications are generated (i.e., $1.0$ correctness score but <0.6<0.6 completeness score)—the clusters at the top of the figure and not overlapping with the gray rectangle. These specifications tend to lead to trivial proofs and hence are filtered out. There is also a big cluster of incorrect specifications that have incorrectly rejected all test cases at the bottom-right corner of the figure (i.e., 0 correctness score but $1.0$ completeness score). There is no way Verus can prove that the Rust function satisfies these specifications, no matter what proof annotations are used, so it is good that SAFE filters all of them out.
@@ E.4 Self-Evolution with Small Backbone Models
Proof Generation Self Debugging (k+kkk+k*k)
@@ F.2 The Effectiveness of Our Self-Debugging Mechanism
Listing 10: An incorrect proof of is_sortedis\_sorted generated by SAFE before self-debugging.
@@
Listing 11: The correct proof of is_sortedis\_sorted fixed by SAFE after self-debugging.
@@ F.3 An Ineffective Example
Listing 12: An incorrect proof of removeremove_oddsodds before and after self-debugging.
@@
Listing 13: The ground truth proof of remove_oddsremove\_odds.
@@ F.4 An Example Proof with a Low-Quality Specification
Listing 14: An example proof of gcdgcd with a low quality specification.